OpenStudy (anonymous):

MEDAL FOR BEST ANSWER The ratio of reds to greens was 5 to 16 and six times the number of reds exceeded the number of greens by 112. How many were reds and how many were greens? Thank you!

OpenStudy (anonymous):

Solve the following for r and g: \[\left\{\frac{r}{g}=\frac{5}{16},6r=g+112\right\} \]

OpenStudy (anonymous):

You should verify yourself that the the following is true: { g = 128, r = 40 }

OpenStudy (anonymous):

Yes, I have the systems, but how do you solve for r and g?

OpenStudy (anonymous):

The second equation is: 6r = g + 112 Solve the above for r.\[r=\frac{g+112}{6} \]Plug r into the first equation and then solve the result for g.

OpenStudy (anonymous):

Okay, thank you so much! :)

OpenStudy (anonymous):

np